Medical cyclotrons need efficient sources of protons to inject into their center. In one kind of ion source, hydrogen atoms (i.e., protons with one orbiting electron) are fed into a chamber where there is a strong magnetic field. Electrons in this chamber are trapped in tight orbits, which greatly increases the chance that they will collide with a hydrogen atom and ionize it. One such source uses a magnetic field of 70 mT, and the electrons' kinetic energy is 1.4 eV. If the electrons travel in a plane perpendicular to the field, what is the radius of their orbits? Express your answer with the appropriate units.
Answer #1

Step 1: Using given kinetic energy, find speed of electron when it enters into magnetic field:

KE = (1/2)*m*V^2

V = sqrt (2*KE/m)

V = sqrt (2*1.4*1.6*10^-19/(9.1*10^-31)) = 701646.415 m/s

Now Using force balance on electron in magnetic field

Fc = Fm

m*V^2/R = q*V*B

R = radius of path = m*V/(q*B)

Using given values

R = 9.1*10^-31*701646.415/(1.6*10^-19*70*10^-3)

R = 57.0*10^-6 m

Let me know if you've any query.